# General Chen-Ricci inequalities for Riemannian submersions and Riemannian maps

### v1

@@ Abstract
We derive general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. We also derive general forms of the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps between Riemannian manifolds, involving the relations between curvatures of subspaces of source and target spaces. These general forms give new techniques that are easy, elegant, and fruitful to obtain the Chen-Ricci inequalities for such smooth mappings with various space forms. As applications, utilizing these general forms, we explicitly establish Chen-Ricci inequalities when the source manifolds of Riemannian submersions and the target manifolds of Riemannian maps belong to broader classes, such as generalized complex and generalized Sasakian space forms, particularly including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and $C(\alpha)$ space forms. Furthermore, we provide Chen-Ricci inequalities for various types of geometrically structured Riemannian submersions and Riemannian maps, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant. In particular, imposing suitable circumstances towards validation, our inequalities become the same as various known inequalities established by many authors in [1, 2, 6, 7, 9, 10, 22, 23, 29, 48, 49].
Keywords: Riemannian submersions; Riemannian maps; Ricci curvatures, Chen-Ricci and improved Chen-Ricci inequalities; space forms; complex and contact manifolds.
@@ 1 Introduction
In submanifold theory, one of the most fundamental problems is to establish simple and sharp relationships between the intrinsic and extrinsic curvature invariants of a Riemannian submanifold [14]. Toward this, a fundamental relation was derived between the Ricci curvature (intrinsic invariant) and the squared mean curvature (extrinsic invariant) for submanifolds of real space forms [12]. Further, such a relation (known as the Chen-Ricci inequality) has also been established for submanifolds of complex, generalized complex, Sasakian, cosymplectic, Kenmotsu, and generalized Sasakian space forms as listed in Table 1. In addition, in [45], Oprea improved the Chen-Ricci inequality for Lagrangian submanifolds of complex space forms, without determining the equality case, using optimization techniques. In [17], Deng also improved this inequality along with the equality case by using some algebraic inequalities. However, improved Chen-Ricci inequalities for Kählerian slant submanifolds of complex space forms and for $C$-totally real submanifolds of Sasakian space forms cannot be obtained using the approaches of Oprea and Deng. Notably, for such cases, Tripathi established an improved Chen-Ricci inequality with equality in [61]. Since then, many authors have established such inequalities for different types of submanifolds in various kinds of ambient manifolds [34, 30, 33, 35, 39, 40, 41, 46, 61, 67]. For a detailed survey on the Chen-Ricci inequalities, we refer to [15] and the references therein.
Table 1: Chen-Ricci inequalities for submanifolds
Type(s) of Submanifolds Ambient Space Form Ref.
totally real, Lagrangian complex [13]
slant, anti-invariant [37]
slant, invariant, totally real generalized complex [24, 58]
(anti-)invariant, slant Sasakian [26]
$C$-totally real [31, 36, 38]
slant, (anti-)invariant [62]
(anti-)invariant, (semi-)slant [16]
(anti-)invariant Kenmotsu [8]
(anti-)invariant, (semi-)slant cosymplectic [66]
semi-slant, (anti-)invariant [32, 66]
semi-invariant [64]
(anti-)invariant, slant generalized Sasakian [25]
slant [57]
The Chen-Ricci inequalities have also been extensively studied by many authors in the context of geometrically structured Riemannian submersions from such kinds of space forms as listed in Table 2. Motivated by the diverse applications of Riemannian submersions mentioned in [19, 51] and the studies mentioned in Table 2, we introduce a general framework of Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. This framework is suitable to derive Chen-Ricci inequalities for Riemannian submersions when the source manifolds are generalized complex and generalized Sasakian space forms, as well as real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, $C(\alpha)$ space forms, etc. In addition, we can use this general framework to establish Chen-Ricci inequalities for various geometrically structured Riemannian submersions, including invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant submersions having the above-mentioned space forms as the source spaces.
@@
The concept of a Riemannian map was introduced by Fischer as a generalization of an isometric immersion and a Riemannian submersion, providing frameworks for quantum models and to compare the geometric properties of two arbitrary Riemannian manifolds [20]. Furthermore, Riemannian maps have been explored in conjunction with complex and contact structures. For detailed studies with applications and examples, see [51] and the references therein. Since Riemannian maps help establish relations between curvature invariants of submanifolds, Lee et al. [29] obtained Chen-Ricci and improved Chen-Ricci inequalities (using Deng’s approach) for Riemannian maps with target spaces as real and complex space forms. Here, we present general forms of the Chen-Ricci and improved Chen-Ricci inequalities (using Tripathi’s approach), along with their equality cases, for Riemannian maps. Notably, these general forms yield the relations between curvatures of subspaces of source and target spaces.
This paper is organized as follows. In Section 2, we provide the definitions of the space forms used in the subsequent sections. Section 3 is dedicated to the Chen-Ricci inequalities for Riemannian submersions. In Subsection 3.1, we provide the necessary preliminaries and foundational concepts required for the subsequent subsections. Subsection 3.2 presents the general forms of Chen-Ricci inequalities for vertical and horizontal distributions. Subsection 3.3 is dedicated to some applications of these general forms. Precisely, we apply these general forms to generalized complex and generalized Sasakian space forms and derive the corresponding Chen-Ricci inequalities, including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and $C(\alpha)$ space forms. In addition, we also obtain and validate the inequalities for various types of Riemannian submersions, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant submersions from these spaces. Remarks 3.7 and 3.11 validate the inequalities obtained with the existing ones. Further, Section 4 is dedicated to the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps. In Subsection 4.1, we give some basic notions required for the studies of subsequent subsections. In Subsection 4.2 (resp. Subsection 4.4), we discuss the general form of the Chen-Ricci (resp. improved Chen-Ricci) inequality. Precisely, Theorems 4.1 and 4.7 are devoted to the main results, that is, the general forms of the Chen-Ricci inequalities and the improved Chen-Ricci inequalities for Riemannian maps. Subsections 4.3 and 4.5 are dedicated to some applications of these general forms. Precisely, we focus on the cases when the target spaces of the Riemannian maps are generalized complex and generalized Sasakian space forms. They also cover special cases when the target spaces are real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and almost $C(\alpha)$ space forms. Similarly to the different types of submanifolds, in Theorems 4.5 and 4.10, we also obtain the inequalities for various kinds of Riemannian maps, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant to these space forms. Remarks 4.4, 4.6, and 4.11 validate the inequalities obtained with the existing ones. In Section 5, we conclude the paper.
@@ 2 Generalized space forms
Let (M,g,J)({M},g,J) be an even-dimensional Riemannian manifold equipped with an almost complex structure $J$ satisfying $J^{2}=-I$, where $I$ is the identity map. Then (M,g,J)({M},g,J) is called an almost Hermitian manifold if for all ${\cal Y}_{1},{\cal Y}_{2}\in\Gamma(T{M})$, the compatibility condition ${g}(J{\cal Y}_{1},J{\cal Y}_{2})=g({\cal Y}_{1},{\cal Y}_{2})$ holds. Additionally, the manifold is referred to as a generalized complex space form, denoted by ${M}(c_{1},c_{2})$, if there exist smooth functions $c_{1}$ and $c_{2}$ on M{M} such that its curvature tensor $R^{M}$ satisfies [44, 60, 65],
@@
$\quad+c_{2}\{g({\cal Y}_{1},J{\cal Y}_{3})J{\cal Y}_{2}-g({\cal Y}_{2},J{\cal Y}_{3})J{\cal Y}_{1}+2g({\cal Y}_{1},J{\cal Y}_{2})J{\cal Y}_{3}\},$
for all ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$. By [63, 65, 60], this includes some special space forms with sectional curvature $c$ for particular values of functions $c_{1}$ and $c_{2}$ as listed in Table 3.
Table 3: Special cases of generalized complex space forms
space form $c_{1}$ $c_{2}$
real $c$ 0
complex $\frac{c}{4}$ $\frac{c}{4}$
real Kähler $\frac{c+3\alpha}{4}$ $\frac{c-\alpha}{4}$
@@
Let $({M},g,\phi,\xi,\eta)$ be an odd-dimensional Riemannian manifold equipped with a $(1,1)$ tensor field $\phi$, a vector field $\xi$, and a $1$-form $\eta$. The structure $({M},g,\phi,\xi,\eta)$ is said to be an almost contact metric manifold if the following conditions are satisfied for all ${\cal Y},{\cal Y}_{1},{\cal Y}_{2}\in\Gamma(T{M})$:
$\eta(\xi)=1,~\phi^{2}({\cal Y})=-{\cal Y}+\eta({\cal Y})\xi~{\rm and~}{g}(\phi{\cal Y}_{1},\phi{\cal Y}_{2})=g({\cal Y}_{1},{\cal Y}_{2})-\eta({\cal Y}_{1})\eta({\cal Y}_{2}).$
Additionally, it holds that $\phi\xi=0$ and $\eta\circ\phi=0$. Furthermore, such a manifold is termed as generalized Sasakian space form (denoted by $M(c_{1},c_{2},c_{3})$) if there exist smooth functions $c_{1}$, $c_{2}$, and $c_{3}$ on $M$ such that the curvature tensor $R^{M}$ of $M$ satisfies [5, 11],
@@
for any ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$. By [5, 11], this also includes several special space forms with sectional curvature $c$ for particular values of functions $c_{1}$ and $c_{2}=c_{3}$ as listed in Table 4.
Table 4: Special cases of generalized Sasakian space forms
space form $c_{1}$ $c_{2}=c_{3}$
Sasakian $\frac{c+3}{4}$ $\frac{c-1}{4}$
Kenmotsu $\frac{c-3}{4}$ $\frac{c+1}{4}$
cosymplectic $\frac{c}{4}$ $\frac{c}{4}$
almost $C(\alpha)$ $\frac{c+3\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$
@@ 3.1 Background
Let $\left(M_{1}^{m_{1}},g_{1}\right)$ and $\left(M_{2}^{m_{2}},g_{2}\right)$ be two Riemannian manifolds. A surjective smooth map $F:\left(M_{1}^{m_{1}},g_{1}\right)$ $\to\left(M_{2}^{m_{2}},g_{2}\right)$ is called a Riemannian submersion if its differential map (i.e. $F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$) is surjective at all $p\in M_{1}$ and $F_{\ast p}$ preserves the length of all horizontal vectors at $p$. Here, the tangent vectors to the fibers (fiber means $\{F^{-1}(q)$ for $q\in M_{2}$}) are referred to as vertical, while the orthogonal vectors to the fibers are termed horizontal. This induces a decomposition of the tangent bundle $TM_{1}$ into a direct sum of two distributions: the vertical distribution $\mathcal{V}=\ker F_{\ast}$ and its orthogonal complement (known as the horizontal distribution) $\mathcal{H}=\left(\ker F_{\ast}\right)^{\perp}$.
@@
3.2 General framework of Chen-Ricci inequalities
In this subsection, we present general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. Toward this, we have three scenarios depending on the horizontal and vertical distributions. We proceed one by one in the following subsubsections.
3.2.1 For vertical distribution
In this subsubsection, we only consider the vertical distribution. In this case, we have the following result.
@@ 3.2.1 For vertical distribution
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})-\frac{1}{4}n^{2}\left\|H\right\|^{2}.$ (3.11)
@@
Since $2\tau_{{\cal V}}^{\ker F_{\ast}}=2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})+2\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}(\vee_{1},\vee_{j},\vee_{j},\vee_{1})$, the equation (3.2.1) implies (3.11). In addition, the equality case is straightforward by (3.12). ∎
3.2.2 For horizontal distribution
In this subsubsection, we only consider the horizontal distribution. In this case, we have the following result.
Theorem 3.2.
${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1}).$ (3.15)
[…]
Proof.
Utilizing (3.3) and (3.6), we obtain
@@ 3.2.2 For horizontal distribution
Substituting (3.2.2) into (3.2.2), we deduce that
@@
This finally gives the desired inequality and the equality case. ∎
3.2.3 For vertical and horizontal distributions
In this subsubsection, we consider both the vertical and horizontal distributions. In this case, we have the following result.
Theorem 3.3.
@@ 3.2.3 For vertical and horizontal distributions
Proof.
By [9, Equation (4.11)], we have the scalar curvature $\tau^{M_{1}}$ of $M_{1}$
@@
$\quad-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}$
@@
$-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
@@
$={\rm Ric}^{\ker F_{\ast}}_{\cal V}\left(\vee_{1}\right)+{\rm Ric}^{\left(\ker F_{\ast}\right)^{\perp}}_{\cal H}\left(h_{1}\right)-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]$
@@
This implies the desired inequality and the equality case. ∎

### v2

@@ Abstract
In this paper, we derive general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. We also derive general forms of the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps between Riemannian manifolds, involving relations between the curvatures of subspaces of the source and target spaces. Further, we illustrate equality cases for all these general forms with two examples. These general forms yield new, easy, and elegant techniques that are fruitful in obtaining the Chen-Ricci inequalities for such smooth mappings with various structured manifolds. As applications, utilizing these general forms, we explicitly establish Chen-Ricci inequalities when the source manifolds of Riemannian submersions and the target manifolds of Riemannian maps belong to broader classes, such as generalized complex and generalized Sasakian space forms, particularly including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and almost $C(\alpha)$ space forms. We also validate our approach by imposing appropriate conditions toward various particular existing cases.
Keywords: Riemannian submersions; Riemannian maps; (improved) Chen-Ricci inequalities; complex and contact manifolds; space forms.
@@ 1 Introduction
In submanifold or immersion theory, one of the most fundamental problems is to establish simple and sharp relationships between the intrinsic and extrinsic curvature invariants of a Riemannian submanifold [11]. Toward this, various authors have derived Chen-Ricci inequalities providing relations between the Ricci curvature (intrinsic invariant) and the squared mean curvature (extrinsic invariant) for different types of submanifolds in several space forms. These inequalities were also improved by Oprea [26] and Deng [13], using optimization techniques and some algebraic inequalities, respectively. However, improved Chen-Ricci inequalities for Kählerian slant submanifolds of complex space forms and for $C$-totally real submanifolds of Sasakian space forms cannot be obtained using the approaches of Oprea and Deng. To avoid such limitations, Tripathi [40] established an approach to improve Chen-Ricci inequalities. For a detailed survey on Chen-Ricci inequalities, we refer to [12] and the references therein.
We know that a Riemannian submersion is the dual concept of isometric immersion and has various applications [15, 32]. The concept of a Riemannian map encompasses the concepts of isometric immersion and Riemannian submersion, providing frameworks for quantum models and for comparing the geometric properties of two arbitrary Riemannian manifolds [16, 32]. Since Riemannian submersions and Riemannian maps help to establish relations between curvature invariants of submanifolds of the source and target manifolds, the Chen-Ricci inequalities have also been studied by many authors in the context of these smooth mappings. For geometrically structured Riemannian submersions, they are explored from various space forms, as listed in Table 1. In addition, for Riemannian maps with real and complex space forms as the target spaces, Lee et al. [22] obtained Chen-Ricci and improved Chen-Ricci inequalities (using Deng’s approach).
@@
Motivatingly, in the current paper, we introduce a general framework of Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds (Theorem 3.1). We also introduce general frameworks of the Chen-Ricci and improved Chen-Ricci inequalities (using Tripathi’s approach) for Riemannian maps between Riemannian manifolds (Theorems 4.1 and 4.2). Notably, these frameworks unify existing studies and yield relations between the curvatures of subspaces of source and target spaces. These frameworks are useful for deriving Chen-Ricci inequalities for Riemannian submersions from arbitrary source manifolds and for Riemannian maps to arbitrary target manifolds (for instance, see Theorems 3.3, 4.4 and 4.5). In addition, we can use these general frameworks to establish Chen-Ricci inequalities for various geometrically structured Riemannian submersions and Riemannian maps, such as invariant, anti-invariant, Lagrangian, semi-invariant, slant, semi-slant, hemi-slant, etc. (see Section 5).
@@ 2 Generalized space forms
Let (M,g,J)({M},g,J) be an even-dimensional almost Hermitian manifold. It is referred to as a generalized complex space form, denoted by ${M}(c_{1},c_{2})$, if its curvature tensor $R^{M}$ satisfies [25, 39, 42],
@@
$\quad+c_{2}\{g({\cal Y}_{1},J{\cal Y}_{3})J{\cal Y}_{2}-g({\cal Y}_{2},J{\cal Y}_{3})J{\cal Y}_{1}+2g({\cal Y}_{1},J{\cal Y}_{2})J{\cal Y}_{3}\}$
for all ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$. Here, $c_{1}$ and $c_{2}$ are smooth functions on M{M}.
@@
Let $({M},g,\phi,\xi,\eta)$ be an odd-dimensional almost contact metric manifold. It is termed as generalized Sasakian space form (denoted by $M(c_{1},c_{2},c_{3})$) if there exist smooth functions $c_{1}$, $c_{2}$, and $c_{3}$ on $M$ such that the curvature tensor $R^{M}$ of $M$ satisfies [5, 10],
@@
for any ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$.
By [5, 10, 41, 42, 39], these generalized space forms include several special space forms based on the particular values of functions $c_{1}$, $c_{2}$, and $c_{3}$ as follows.
Table 2: Special cases
space form $c_{1}$ $c_{2}$ $c_{3}$
real $c$ 0 -
complex $\frac{c}{4}$ $\frac{c}{4}$ -
real Kähler $\frac{c+3\alpha}{4}$ $\frac{c-\alpha}{4}$ -
Sasakian $\frac{c+3}{4}$ $\frac{c-1}{4}$ $\frac{c-1}{4}$
Kenmotsu $\frac{c-3}{4}$ $\frac{c+1}{4}$ $\frac{c+1}{4}$
cosymplectic $\frac{c}{4}$ $\frac{c}{4}$ $\frac{c}{4}$
almost $C(\alpha)$ $\frac{c+3\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$
@@ 3.1 Background
Let $\left(M_{1}^{m_{1}},g_{1}\right)$ and $\left(M_{2}^{m_{2}},g_{2}\right)$ be two Riemannian manifolds. A smooth surjective map $F:\left(M_{1}^{m_{1}},g_{1}\right)$ $\to\left(M_{2}^{m_{2}},g_{2}\right)$ is called a Riemannian submersion if its differential map (i.e., $F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$) is surjective at all $p\in M_{1}$ and $F_{\ast p}$ preserves the length of all horizontal vectors at $p$. Here, the tangent vectors to the fibers (fiber means $\{F^{-1}(q)$ for $q\in M_{2}$}) are termed vertical, while the orthogonal vectors to the fibers are termed horizontal. This induces a decomposition of the tangent bundle $TM_{1}$ into a direct sum of two distributions: the vertical distribution $\mathcal{V}=\ker F_{\ast}$ and its orthogonal complement (known as the horizontal distribution) $\mathcal{H}=\left(\ker F_{\ast}\right)^{\perp}$.
@@
3.2 General Chen-Ricci inequalities
Here, we present general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. Toward this, we have the following result.
@@ 3.2 General Chen-Ricci inequalities
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.11)
${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1}),$ (3.12)
and
Further, equality cases are as follows.
  1. $(1)$
@@
[…]
    1. $(a)$
      totally geodesic when $n>2$;
    2. $(b)$
      totally umbilical when $n=2$.
  2. $(2)$
[…]
@@ 2. $(2)$
Since $2\tau_{{\cal V}}^{\ker F_{\ast}}=2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})+2\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}(\vee_{1},\vee_{j},\vee_{j},\vee_{1})$, the equation (3.2) implies (3.11). In addition, its equality case is straightforward by (3.14).
Further, utilizing (3.3) and (3.6), we obtain
@@
Substituting (3.2) into (3.2), we deduce that
@@
Furthermore, by [8, Equation (4.11)], we have the scalar curvature $\tau^{M_{1}}$ of $M_{1}$
@@
$\quad-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]^{2}-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}$
@@
$-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]^{2}-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
@@
$={\rm Ric}^{\ker F_{\ast}}_{\cal V}\left(\vee_{1}\right)+{\rm Ric}^{\left(\ker F_{\ast}\right)^{\perp}}_{\cal H}\left(h_{1}\right)-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]^{2}$
@@
[…]
3.3 Example
Here, we provide an example satisfying the equality of Chen-Ricci inequalities derived in Theorem 3.1, as follows.
Example 3.2.
Considering $\left(M_{1}=(-\frac{\pi}{2},\frac{\pi}{2})\times_{\cos t}\mathbb{C},~g_{1}=dt^{2}+\cos^{2}t\sum\limits_{i=1}^{3}((dx_{i})^{2}+(dy_{i})^{2})\right)$ and $\left(M_{2}=\mathbb{R}^{4},g_{2}=\cos^{2}t\sum\limits_{j=1}^{4}(dz_{j})^{2}\right)$, we define $F:(M_{1},g_{1})\to(M_{2},g_{2})$ such that
$F(t,x_{1},y_{1},x_{2},y_{2},x_{3},y_{3})=(x_{2},y_{2},x_{3},y_{3}).$
Then we get
$(\ker F_{\ast})=\operatorname{span}\left\{\vee_{1}=\frac{\partial}{\partial t},\vee_{2}=\frac{1}{\cos t}\frac{\partial}{\partial x_{1}},\vee_{3}=\frac{1}{\cos t}\frac{\partial}{\partial y_{1}}\right\},$
$(\ker F_{\ast})^{\perp}=\operatorname{span}\left\{h_{1}=\frac{1}{\cos t}\frac{\partial}{\partial x_{2}},h_{2}=\frac{1}{\cos t}\frac{\partial}{\partial y_{2}},h_{3}=\frac{1}{\cos t}\frac{\partial}{\partial x_{3}},h_{4}=\frac{1}{\cos t}\frac{\partial}{\partial y_{3}}\right\},$
and
$({\rm range}~F_{\ast})=\operatorname{span}\left\{F_{\ast}(h_{j})=\frac{1}{\cos t}\frac{\partial}{\partial z_{j}}\right\}_{j=1}^{4},$
where $\left\{\frac{\partial}{\partial t},\frac{1}{\cos t}\frac{\partial}{\partial x_{1}},\frac{1}{\cos t}\frac{\partial}{\partial y_{1}},\frac{1}{\cos t}\frac{\partial}{\partial x_{2}},\frac{1}{\cos t}\frac{\partial}{\partial y_{2}},\frac{1}{\cos t}\frac{\partial}{\partial x_{3}},\frac{1}{\cos t}\frac{\partial}{\partial y_{3}}\right\}$ and $\left\{\frac{\partial}{\partial z_{j}}\right\}_{j=1}^{4}$ are bases of $T_{p}M_{1}$ and $T_{F(p)}M_{2}$, respectively. For all $i,j\in\{1,2,3,4\}$, we verify that
$g_{1}(h_{i},h_{j})=g_{2}(F_{\ast}h_{i},F_{\ast}h_{j}).$
Therefore, $F$ is a Riemannian submersion between Riemannian manifolds.
We calculate the non-vanishing Christoffel symbols for the metric $g_{1}$, which are $\Gamma^{t}_{i,i}=\sin t\cos t$ and $\Gamma^{i}_{t,i}=-\tan t$ for all $i\in\{x_{j},y_{j}\}_{j=1}^{3}$. Using these values, we calculate the following:
$\nabla^{M_{1}}_{\vee_{1}}\vee_{i}=\nabla^{M_{1}}_{\vee_{2}}\vee_{3}=\nabla^{M_{1}}_{\vee_{3}}\vee_{2}=0~\text{for all $1\leq i\leq 3$},$
$\nabla^{M_{1}}_{\vee_{2}}\vee_{1}=-\tan t\sec t\frac{\partial}{\partial x_{1}},\nabla^{M_{1}}_{\vee_{3}}\vee_{1}=-\tan t\sec t\frac{\partial}{\partial y_{1}},\nabla^{M_{1}}_{\vee_{2}}\vee_{2}=\nabla^{M_{1}}_{\vee_{3}}\vee_{3}=\tan t\frac{\partial}{\partial t},$
$\nabla^{M_{1}}_{h_{i}}h_{j}=0~\text{for all $1\leq i\neq j\leq 4$ and ~}\nabla^{M_{1}}_{h_{i}}h_{i}=\tan t\frac{\partial}{\partial t}~\text{for all $1\leq i\leq 4$}.$
Thus, $\mathcal{H}\nabla^{M_{1}}_{\vee_{i}}\vee_{j}=T_{\vee_{i}}\vee_{j}=0$ for all $1\leq i,j\leq 3$. Hence, for all $h_{t}\in\{h_{1},h_{2},h_{3},h_{4}\}$, we have
$T_{ij}^{t}=g_{1}(T_{\vee_{i}}\vee_{j},h_{t})=0,\quad 1\leq i,j\leq 3~\text{and ~}1\leq t\leq 4.$
[…]